# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### 27 Feb 2024 v1

@@ 5 Numerical experiments
  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ has better clustering properties than the baseline preconditioned matrices, consequently improves the computational performance of our proposed PESS PGMRES process. Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.08330]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ they are in $(0,1.9983],$ which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively. For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3.
@@
  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},$ and ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ and $9.5124e+09,$ respectively, which are very large. Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.422$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
@@
The above discussions affirm that the proposed PESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
@@
$\Delta B=\varepsilon*N_{P}*std(B).*randn(m,n)\quad\text{and}\quad\Delta C=\varepsilon*N_{P}*std(C).*randn(p,m),$
where $\varepsilon=10^{-4}$ and $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively. Here, $randn(m,n)$ is the is an $m$-by-$n$ normally distributed random matrix, generated by the MATLAB command $randn$. We perform the numerical test for $l=16,32,48,64$ and $80$ for Example 1 using the PESS PGMRES process (Case II, $s=12$). The calculated norm error $\|\tilde{\textbf{u}}-\textbf{u}\|_{2}$ among the solution of the perturbed system and the original system with increasing $N_{P}$ from $5\%$ to $40\%$ with step size $5\%$ are displayed in the figure 4. We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is not sensitive to small perturbation on $\mathcal{A}$.
Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, and PESS PGMRES processes for Example 2.
@@
Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, and PESS preconditioners in Case II for Example 2.
@@
  • In Case I: $\alpha=0.02$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.02,\beta=0.2,\gamma=0.002$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha I,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma I$ for ${\mathscr{P}}_{PESS}.$
  • In Case II: $\alpha=2$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=2,\beta=2,\gamma=0.002$ and $P=A,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=\frac{1}{2}\alpha A,\Lambda_{2}=\frac{1}{2}\beta I,\Lambda_{3}=\frac{1}{2}\gamma CC^{T}.$
For the IBD preconditioner, the matrices $\hat{A}$ and $\hat{S}$ are constructed as in Example 1. In the interval $[10,30]$, the empirical optimal choice for $s$ is found to be $25$ for Case I and $26$ for Case II. Table 3 shows that the GMRES process has a very slow convergence speed and also does not converge for $\textbf{h}=1/64$ within $7000$ iterations. The proposed PESS preconditioner requires almost five times fewer iterations compared to the IBD preconditioner for convergence. Moreover, in both cases, the PESS preconditioned process outperforms the SS, RSS, and EGSS preconditioners in terms of IT and CPU times. Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS, RSS, and EGSS, IT increases as the size of $\mathcal{A}$ increases. Furthermore, in Case II with $l=1/64$, our proposed PESS preconditioner is approximately $18\%$, $91\%$, $60\%$, and $19\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are observed for other values of $l.$
The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS preconditioner compared to the baseline IBD, SS, RSS, EGSS, and PESS preconditioners in terms of RES versus IT counts.
@@
Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
(a) CPU versus parameters $s$ and $\alpha$
(b) IT versus parameters $s$ and $\alpha$
Figure 7: Characteristic curves for CPU (left) and IT (right) of the proposed PESS iteration process by varying $s$ in the interval $[1,30]$ with step size $1$ and $\alpha$ in the interval $[0.05,1]$ with step size $0.05$ with $\textbf{h}=1/32$ in Case I for Example 2.
Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS PGMRES process. Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.07e+05.$ While for Case I and $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.4024,$ indicating that the proposed PESS preconditioned system is well-conditioned, ensuring an efficient and robust solution. In addition, to demonstrate the relationship of the parameters and the convergence speed of the PESS preconditioner, we plot the graph of CPU times and IT counts by varying the parameters $s$ and $\alpha,$ in Figure 7. For $\textbf{h}=1/32$, $s$ varying in the interval $[1,30]$ with step size one and $\alpha$ from the interval $[0.05,1]$ with step size $0.05$ Other parameters remain consistent with Case I. We draw the following observations from Figure 7.
  • For the given value of $s$, the proposed PESS preconditioner is insensitive to the change of the parameter $\alpha.$
  • With an increase in $s$, IT and CPU times show a substantial decrease irrespective of the value of $\alpha.$
@@ 6 Conclusion
In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ to solve the three-by-three block SPP (1.1). For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived. Moreover, we estimated the spectral bounds of the proposed PESS preconditioned matrices. This empowers us to derive spectral bounds for SS and EGSS preconditioned matrices, which have not yet been investigated in the literature to the best of our knowledge. Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS preconditioner. The key observations are as follows: $(i)$ the proposed PESS preconditioner is found to outperform the existing baseline preconditioners in terms of IT and CPU times. $(ii)$ The proposed PESS preconditioner significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs. $(iii)$ The PESS preconditioned matrix has a more clustered spectral distribution than the baseline preconditioned matrices. $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner.
Acknowledgements
The second author (Pinki Khatun) sincerely acknowledges the Council of Scientific &\&& Industrial Research (CSIR) in New Delhi, India, for financial assistance in the form of a fellowship (File no. 09/1022(0098)/2020-EMR-I).
@@ Acknowledgements
Data availability
No data was used for the research described in the article.

### 17 May 2024 v2

@@ 6 Numerical experiments
  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LESS}^{-1}\mathcal{A}$ have better clustering properties than the baseline preconditioned matrices, consequently improves the computational efficiency of our proposed PESS and LPESS PGMRES processes. Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.0714]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ are in $(0,1.9991]$ and $(0,1],$ respectively, which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively. For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3.
@@
  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},$ and ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ and $9.5124e+09,$ respectively, which are very large. Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.4221$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
@@
The above discussions affirm that the proposed PESS and LPESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
@@
$\Delta B=10^{-4}*N_{P}*std(B).*randn(m,n)\quad\text{and}\quad\Delta C=10^{-4}*N_{P}*std(C).*randn(p,m),$
where $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively. Here, $randn(m,n)$ is the is an $m$-by-$n$ normally distributed random matrix, generated by the MATLAB command $randn$. We perform the numerical test for $l=16,32,48,64$ and $80$ for Example 1 using the PESS PGMRES process (Case II, $s=12$). The calculated norm error $\|\tilde{\textbf{u}}-\textbf{u}\|_{2}$ among the solution of the perturbed system and the original system with increasing $N_{P}$ from $5\%$ to $40\%$ with step size $5\%$ are displayed in the figure 4. We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is not sensitive to small perturbation on $\mathcal{A}$.
Example 2.
[…]
$-\Delta{\bf u}+\nabla p=0,\quad\text{in}\quad\Xi,$ (6.1)
$\quad\quad\quad\nabla\cdot{\bf u}=0,\quad\text{in}\quad\Xi.$
[…]
Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, PESS and LPESS PGMRES processes for Example 2.
@@
Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, PESS and LPESS preconditioners in Case II for Example 2.
@@
Table 3 shows that the GMRES process has a very slow convergence speed and also does not converge for $\textbf{h}=1/64$ within $7000$ iterations. The proposed preconditioners require almost five times fewer iterations compared to the IBD preconditioner for convergence. Moreover, in both cases, the PESS and LPESS preconditioners outperform the SS, RSS, and EGSS preconditioners in terms of IT and CPU times. Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS and EGSS, IT increases as the size of $\mathcal{A}$ increases. Furthermore, in Case II with $l=1/64$, our proposed LPESS preconditioner is approximately $28\%$, $90\%$, $53\%$, and $16\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are observed for other values of $l.$
@@
(g) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS and LPESS PGMRES processes compared to the baseline IBD, SS, RSS, and EGSS ones in terms of RES versus IT counts. Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS and LPESS PGMRES processes.
(a) PESS PGMRES process
(b) LPESS PGMRES process
Figure 7: Characteristic curves for IT of the proposed PESS (left) and LPESS (right) PGMRES processes by varying $s$ in the interval $[1,100]$ with step size $1$ with $\textbf{h}=1/16$ in Case I for Example 2.
Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.0701e+05.$ While for Case I with $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.3353$ and $\kappa({\mathscr{P}}_{LPESS}^{-1}\mathcal{A})=1.2056,$ indicating that the proposed PESS and LPESS preconditioned systems are well-conditioned, ensuring an efficient and robust solution. In addition, to demonstrate the relationship of the parameter $s$ and the convergence speed of the PESS and LPESS preconditioners, we plot graphs of IT counts by varying the parameters $s$ in the interval $[1,100]$ with step size one in Figure 7. We consider $\textbf{h}=1/16$ and other choices for $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ as in Case I. Figure 7 shows that, with the increasing value of $s,$ decreasing trend in the $IT$ counts for both the PESS and LPESS preconditioner. Moreover, for $s>22,$ using LPESS preconditioner, we can solve the three-by-three $SPP$ mentioned only in $3$ iterations.
@@ 7 Conclusion
In this paper, we proposed the PESS iterative process and corresponding PESS preconditioner ${\mathscr{P}}_{PESS}$ and its relaxed variant LPESS preconditioner to solve the three-by-three block SPP (1.1). For the convergence of the proposed PESS iterative process, necessary and sufficient criteria are derived. Moreover, we estimated the spectral bounds of the proposed PESS and LPESS preconditioned matrices. This empowers us to derive spectral bounds for SS and EGSS preconditioned matrices, which have not yet been investigated in the literature to the best of our knowledge. Numerous experimental analyses are performed to demonstrate the effectiveness of our proposed PESS and LPESS preconditioners. The key observations are as follows: $(i)$ the proposed PESS and LPESS preconditioners are found to outperform the existing baseline preconditioners in terms of IT and CPU times. $(ii)$ The proposed preconditioners significantly reduces the condition number of $\mathcal{A}$, consequently showing the its proficiency in solving three-by-three block SPPs. $(iii)$ The PESS and LPESS preconditioned matrices have a more clustered spectral distribution than the baseline preconditioned matrices. $(iv)$ Sensitivity analysis conducted by introducing different percentages of noise on the system (1.1) showcases the robustness of the proposed PESS preconditioner. Further investigation to determine the optimal parameter $s$ can be done in future research endeavors.
@@
Acknowledgments
Pinki Khatun sincerely acknowledges the Council of Scientific &\&& Industrial Research (CSIR) in New Delhi, India, for financial assistance in the form of a fellowship (File no. 09/1022(0098)/2020-EMR-I).